OpenStudy (anonymous):

The cost of eleven CDs and one DVDs is $69.41; and the cost of seven DVDs and twelve CDs is $112.11. How much do a CD and a DVD cost?

OpenStudy (anonymous):

x=5.75 y=6.16

OpenStudy (anonymous):

11x + y =69.41 12x+7y=112.11 12x/7+1y=112.11/7 11x-12x/7 = 69.41-112.11/7 65/7x = 373.76/7 65x=373.76 x=373.76/65 x= 5.75 then work in x for y